OpenStudy (anonymous):

Which equation does NOT show direct variation? 12y = x y = ½x y + 20 = x + 20 y = 34x + 2

OpenStudy (phi):

direct variation means you can write the equation as y = k x where k is a number (such as a fraction, or integer, etc) for the first equation 12y= x if you divide both sides by 12 you will get \[ y = \frac{1}{12} x\] the second equation is \[ y= \frac{1}{2}x\] the 3rd equation y+20= x+20 if you add -20 to both sides you get y+20-20 = x+20-20 or y+0=x+0 or y= x which can be written as \[ y = 1 x \]